Let the plane P pass through the 3 points R = (1, 1, 6), S= (2,5,4) and T = (1,2,3).


Find the values of a, b, c and d in the equation "ax + by + cz = d" of plane P.



1
Expert's answer
2021-07-01T06:14:13-0400

The plane through three points


"\begin{vmatrix}
 x-x_1 & y-y_1 & z-z_1 \\
 x_2-x_1 & y_2-y_1 & z_2-z_1 \\
 x_3-x_1 & y_3-y_1 & z_3-z_1 \\
\end{vmatrix}=0"

"\begin{vmatrix}
 x-1 & y-1 & z-6 \\
 2-1 & 5-1 & 4-6 \\
 1-1 & 2-1 & 3-6 \\
\end{vmatrix}=0"

"(x-1)\begin{vmatrix}
 4 & -2 \\
 1 & -3
\end{vmatrix}-(y-1)\begin{vmatrix}
 1 & -2 \\
 0 & -3
\end{vmatrix}+(z-6)\begin{vmatrix}
 1 & 4 \\
 0 & 1
\end{vmatrix}=0"

"-10(x-1)+3(y-1)+z-6=0"

"10x-3y-z=1"


"a=10, b=-3, c=-1, d=1"